OpenStudy (anonymous):

Need an answer checked! Mai borrowed $12,000 at 5% interest for 6 years. What was the total interest Mai paid? $40000?

OpenStudy (mahoganie.carson):

So if it is simple interest then basically you are looking for the interest paid so 5%=5/100=0.05 assuming that the interest is calculated yearly (12000 times 0.05) times 6 because 6 years so if you calculate that out then you get 3600 Mai paid $3600 total

OpenStudy (anonymous):

Hm, is it not 12,000=p*0.05*6 Did I do something wrong there?